Find the infimum and supremum of {x ∈ R : x2+2x=3}


1
Expert's answer
2021-09-01T11:53:56-0400

"\{x ∈ R : x^2+2x=3\}=\{x ∈ R : (x+1)^2=4\}=\{1, -3\}"

Therefore, "\inf \{x ∈ R : x^2+2x=3\}=\inf\{1, -3\}=-3",

"\sup \{x ∈ R : x^2+2x=3\}=\sup\{1, -3\}=1"